Write y-4=-3/2(x+6) in standare form.
Not in book. Need help please

Answer by stanbon(75887)

Write y-4=-3/2(x+6) in standare form.
y = (-3/2)x-9+4
y=(-3/2)x-5
Multiply thru by 2 to get:
2y=-3x-10
3x+2y=-10
Cheers,
Stan H.